OpenStudy (amy0799):

The function f and g are differentiable at x=10 and x=20 and f(g(x))=x^2. if f(10)=5, f'(10)=4, f'(20)=-5, and g(10)=20, what is the value of g'(10)?

OpenStudy (freckles):

have you used chain rule to differentiate f(g(x)) yet?

OpenStudy (amy0799):

I learned how to do the chain rule

OpenStudy (freckles):

ok then what is \[\frac{d}{dx}f(g(x))=?\]

OpenStudy (amy0799):

f(x)g'(x)+g(x)f'(x)

OpenStudy (freckles):

that is the product rule we don't have f*g we have f composed with g

OpenStudy (freckles):

this is why I asked you to use chain rule to differentiate f(g(x))

OpenStudy (amy0799):

I thought u use the chain rule when its something to a power

OpenStudy (freckles):

power rule is what you use when you have a constant power

OpenStudy (freckles):

chain rule is what you use when you have a function inside a function

OpenStudy (amy0799):

ooh ok. So how would I do the chain rule for this? would it be f'(g(x))*g'(x)?

OpenStudy (freckles):

that is right

OpenStudy (freckles):

\[f(g(x))=x^2 \\ \text{ differentiating both sides } \\ f'(g(x)) \cdot g'(x)=2x\]

OpenStudy (freckles):

not enter in 10 for x

OpenStudy (freckles):

\[f'(g(10)) \cdot g'(10)=2(10)\]

OpenStudy (freckles):

you are given g(10)

OpenStudy (freckles):

g(10)=20 right?

OpenStudy (freckles):

so replace g(10) with 20 \[f'(g(10)) \cdot g'(10)=2(10)\] \[f'(20) \cdot g'(10)=2(10)\]

OpenStudy (freckles):

see if you can finish the rest

OpenStudy (amy0799):

g'(10)=25?

OpenStudy (freckles):

\[f'(20)=-5 \\ \text{ so we have } \\ -5 \cdot g'(10)=2(10)\]

OpenStudy (freckles):

hmmm how did you get g'(10)=25?

OpenStudy (amy0799):

oh I added, oops

OpenStudy (freckles):

you do know the operation between -5 and g'(10) is multiplication and not addition :p

OpenStudy (amy0799):

g'(10)=-4?

OpenStudy (freckles):

20/-5 is -4 good worrk

OpenStudy (amy0799):

thank you!

OpenStudy (amy0799):

I have another question, do u mind helping me till?

OpenStudy (freckles):

I can take a look

OpenStudy (amy0799):

In the table below, the values of f(x), g(x), f'(x) and g'(x) are given 2 values of x. if y =[f(2x)+g(x)]^2, find y'(3)

OpenStudy (freckles):

well we know we are going to have differentiate since we want to find y'(of something)

OpenStudy (amy0799):